4. What is the Molality of a solution containing 7.78g of Urea [(NH2)2CO2] in 203g of
water?
5. Lead is a poisonous metal that especially affects children because they retain a
larger fraction of lead than adults do. Lead levels of 0.250ppm in a child cause delayed
cognitive development. How many moles of lead present in 1.00g of child’s blood
would 0.250ppm represent.
6. Acetone, C3H6, is the main ingredient of nail polish remover. A solution is made up
by adding 35.0mL of acetone (d=0.790g/mL) to 50.0mL of ethyl alcohol, C2H6O
(d=0.789g/mL). Assuming volumes are additive, calculate (a) the mass percent of
acetone in the solution. (b) the volume percent of ethyl alcohol in the solution. (c) the
mole fraction of acetone in the solution.
1
Expert's answer
2021-09-16T02:01:30-0400

Number four

Step by step solution


Molality can be calculated by the following formula.


From the given,


Mass of urea = 7.78 g


Let’s convert the mass of urea into moles.


Molar mass of urea = 60.055 g/mol


Let’s calculate the moles of urea.




Hence, moles of urea are 0.1295.



"Molarity= moles\/litre"


Moles = 7.78/60 =0.1297moles

Molarity= (0.1297× 0.203)/1000

= 2.63×10-3M


Number five








Moles of Lead =2.67moles


Number six

a) 35 mL (0.790 g/mL) = 27.65 g acetone

50 mL (0.789 g/mL) = 39.45 g ethyl alcohol

total = 67.1 g

( 27.65 / 67.1 )100


= 41.2 mass% acetone


b)the volume percent of ethyl alcohol in the solution


(50 / 85 )100


= 58.8 vol% ethyl alcohol


c)Convert grams from part a into moles acetone and moles EtOH.


Sum moles to get a total moles.


Then

mole fraction acetone


= (moles acetone/total moles)

=Mole fraction of ethyl alcohol = moles of ethyl alcohol/total moles


Mole fraction of acetone=0.46